OpenStudy (anonymous):

(1/16x^4+3/4x^3+1/8x^2+6)+(-9/16x^4+7/8x^2-6) please help.

OpenStudy (turingtest):

to confirm the expression is\[(1/16x^4+3/4x^3+1/8x^2+6)+(-9/16x^4+7/8x^2-6)\] ignore the parentheses; they mean nothing here. combine like terms and common denominators \[=(1-9)/16x^4+3/4x^3+(1+7)/16x^2+6-6\]\[=-1/2x^4+3/4x^3+1/x^2\]